# Balancing Efficiency and Expressiveness: Subgraph GNNs with Walk-Based Centrality

### v1

@@ 4.1 Our Approach
$C_{v}^{\mathsf{CSE}}=\left[1,\frac{(A)^{1}_{vv}}{1!},\frac{(A)^{2}_{vv}}{2!},\frac{(A)^{3}_{vv}}{3!},\dotsc,\frac{(A)^{K}_{vv}}{K!}\right]\in\mathbb{R}^{K+1}.$ (4)
As $K\to\infty$, the sum of these terms clearly coincides with the SC of the node for the default $\beta=1$. We refer readers to Section E.2, for considerations on how CSEs compare with RWSEs.
@@
Algorithm 1 Hybrid Marking Network
1:Graph $G=(A,X)$, Subgraph GNN $f$, Max walk-length $K$, Number of marks $T$
2:$C^{\mathsf{CSE}}_{v,k}\leftarrow 0\quad\forall v\in G,k\in[K]$triangleright SE init.
3:$M_{v,t}\leftarrow 0\quad\forall v\in G,t\in[T]$triangleright Mark init.
4:$B\leftarrow I$
5:for $k\in[K]$ do
6:     for $v\in G$ do
7:         $C^{\mathsf{CSE}}_{v,k}\leftarrow\nicefrac{{B_{vv}}}{{k!}}$triangleright Compute SEs
8:     end for
9:     $B\leftarrow B\cdot A$
10:end for
11:$\tilde{C}^{SC}\leftarrow\sum_{k}C^{\mathsf{CSE}}_{:,k}$triangleright Estimate SC
12:$\mathcal{M}\leftarrow\text{select-top}(\tilde{C}^{SC},T)$triangleright Select nodes
13:for $t\in[T]$ do
14:     $M_{\mathcal{M}[t],t}\leftarrow 1$triangleright Mark nodes
15:end for
16:$y_{G}\leftarrow f(A,X\oplus C^{\mathsf{CSE}},M)$triangleright Forward-pass
17:return $y_{G}$
In view of (1) and (2), we dub our approach HyMN, as in Hybrid Marking Network. These steps are depicted in Figure 2 and described in Algorithm 1. We note the following. First, for a large enough $K$, centrality values can be approximated by directly summing over the computed $K$-order CSEs. Second, node marking does not require any alteration of the original graph topology, making it unnecessary to store the subgraph connectivity. We thus opt not to materialize the bag of subgraphs: we only record marking information in the feature tensor and implement a custom message-function that processes it in an equivariant way. From an engineering standpoint, this allows for further memory-complexity enhancement w.r.t. generic Subgraph GNN approaches.
Figure 2: An overview of the pipeline for HyMN. We first calculate the centrality encoding for each node in the graph. We sum each row of this structural encoding to approximate the Subgraph Centrality of each node. We then sample $T$ node-marked subgraphs with the highest Subgraph Centrality on the marked node and concatenate the centrality encoding with the initial node features.
4.2 Expressivity of Subgraph GNNs with Centrality Encodings
@@
We answer these questions with graph separation arguments (Xu et al., 2018; Morris et al., 2019) and highlight how, in fact, the two approaches are not generally comparable. We demonstrate that subsampling Subgraph GNNs with high-centrality marking does not subsume CSE-enhanced MPNNs, while, at the same time, the discriminative power of the former approach is not fully captured by the latter. Since neither technique alone fully subsumes the other, our analysis emphasizes the advantages of combining them in HyMN for improved expressiveness. Proofs and additional details are reported in Appendix C.
@@
This result exposes a limitation of subsampled Subgraph GNNs in distinguishing between two non-isomorphic graphs with differing closed walks, features which are, instead, captured by CSEs. Notably, as discussed in Proposition 2 (Appendix C), a full-bag approach is capable of capturing CSEs. This observation suggests that while CSEs do not universally enhance the expressiveness of any Subgraph GNN, they can be beneficial when subsampling a limited number of subgraphs.
Taken together, Theorems 1 and 2 indicate that leveraging SC both as a structural encoding and as a means for subgraph sampling is advantageous from an expressiveness perspective, justifying the integration of the two techniques in HyMN.
@@ 5 Experiments
  • (Q2) Can HyMN efficiently scale to graphs out of reach for Subgraph GNNs?
@@
Synthetic experiment for counting substructures. The ability of a model to count local substructures is an acknowledged way of evaluating its expressive power (Bouritsas et al., 2022; Arvind et al., 2020; Tahmasebi et al., 2020). In order to answer (Q1) and test the efficacy of subgraph sampling with node centrality, we explored the ability of a Subgraph GNN to count different small substructures as we increase the number of subgraphs in our bag. We closely followed the experimental procedure of (Chen et al., 2020), but modified the data generation process to render the task more challenging and informative [^5] In particular, we considered larger graphs with a similar number of nodes to correct an undesired correlation between the graph size and the task targets observed in the original data (see Section F.2).. We compared the performance of sampling subgraphs based on different approaches: random sampling and sampling based on the highest values of different centrality measures. No CSEs are employed in this setting. Fig. 3 reports results for triangle and $4$-cycle counting. These demonstrate the significant improvement afforded by high-centrality sampling over random sampling with fewer subgraphs (Q1). Additionally, they show that SC-based sampling generally outperformed other centrality measures, demonstrating the benefits of focusing on walk-based centralities, in alignment with our analysis in Section 3. The full set of results for additional substructures is reported in Appendix E.
@@
Figure 3: Comparing Random and Centrality-based sampling on counting different substructures in the graph. For reference, the average graph size is 59.33, corresponding to the average bag-size of a full-bag Subgraph GNN.
OGB. We tested HyMN on several datasets from the OGB benchmark (Hu et al., 2020b). To further examine (Q1), Table 2 shows the performance of our approach in relation to MPNNs, a full-bag Subgraph GNN and other subgraph sampling policies with the same number of subgraphs. Notably, even without using the centrality-based encoding (HyMN w/out CSE), our method matches the performance of a learnable sampling policy (POLICY-LEARN (Bevilacqua et al., 2024)) and consistently outperforms MPNNs and random sampling policies. Additionally, we observe that augmenting node features with CSEs can significantly increase performance on MOLHIV and MOLBACE, outperforming a full-bag Subgraph GNN (Q4). These results suggest that centrality sampling is effective and that additionally incorporating centrality information can lead to performance improvements on real-world datasets, aligning with our findings in Section 3. Additional results for MOLHIV are reported in Table 4, where we show HyMN can outperform strong GNN baselines (Q3). Run-time comparisons for this dataset are additionally enclosed in Appendix G.
@@
Table 2: Results on OGB datasets. The first and second best results for each task are color-coded.
Method MOLHIV (ROC-AUC $\uparrow$) MOLBACE (ROC-AUC $\uparrow$) MOLTOX21 (ROC-AUC $\uparrow$)
GCN Kipf & Welling (2017) 76.06 ±$\pm 0.97$ 79.15 ±$\pm 1.44$ 75.29 ±$\pm 0.69$
GIN Xu et al. (2018) 75.58 ±$\pm 1.40$ 72.97 ±$\pm 4.00$ 74.91 ±$\pm 0.51$
FULL Bevilacqua et al. (2024) 76.54 ±$\pm 1.37$ 78.41 ±$\pm 1.94$ 76.25 ±$\pm 1.12$
OSAN Qian et al. (2022) - 76.30 ±$\pm 3.00$
MAG-GNN Kong et al. (2024) 77.12 ±$\pm 1.13$
RANDOM (T = 2) Bevilacqua et al. (2024) 77.55 ±$\pm 1.24$ 75.36 ±$\pm 4.28$ 76.65 ±$\pm 0.89$
POLICY-LEARN (T = 2) Bevilacqua et al. (2024) 79.13 ±$\pm 0.60$ 78.40 ±$\pm 2.85$ 77.47 ±$\pm 0.82$
HyMN (GIN, T=2) w/out CSE 79.77 ±$\pm 0.70$ 78.22 ±$\pm 4.02$ 77.68 ±$\pm 0.71$
@@
RANDOM (T=5) Bevilacqua et al. (2024) 77.30 ±$\pm 2.56$ 78.14 ±$\pm 2.36$ 76.62 ±$\pm 0.89$
POLICY-LEARN (T=5) Bevilacqua et al. (2024) 78.49 ±$\pm 1.01$ 78.39 ±$\pm 2.28$ 77.36 ±$\pm 0.60$
HyMN (GIN, T=5) w/out CSE 79.62 ±$\pm 1.14$ 78.57 ±$\pm 1.31$ 77.82 ±$\pm 0.59$
@@
Peptides. In order to evaluate the ability of HyMN to scale to larger graphs (Q2), we experimented on the Peptides datasets from the LRGB benchmark (Dwivedi et al., 2022). The average number of nodes in these graphs is 150.94, so it is difficult for a full-bag Subgraph GNN to process. Additionally, using the centrality encoding to sample just one or two additional node-marked subgraphs can improve performance on both datasets. We also outperform GPS (Rampášek et al., 2022) (a Graph Transformer) and match the performance of Graph-MLP-Mixer (He et al., 2023) (Q3). Our timing experiments on Peptides-Func demonstrate that we are significantly more efficient than both of these approaches, only increasing the training time per epoch over a GCN by $42\%$ and the inference time by $25\%$ using ‘HyMN (GCN, T=1)’ (Q2).
Method Precompute (s) Train (s/epoch) Test (s) Peptides-Func ($\uparrow$) Peptides-Struct ($\downarrow$)
@@
GPS (Rampášek et al. (2022)) 20.87 ±$\pm 0.43$ 8.39 ±$\pm 0.05$ 0.611 ±$\pm 0.005$ 0.6535 ±$\pm 0.0041$ 0.2500 ±$\pm 0.0005$
Graph-ViT (He et al. (2023)) 29.12 ±$\pm 0.61$ 6.78 ±$\pm 0.01$ 0.709 ±$\pm 0.009$ 0.6942 ±$\pm 0.0075$ 0.2449 ±$\pm 0.0016$
G-MLP-Mixer (He et al. (2023)) 29.52 ±$\pm 0.69$ 6.87 ±$\pm 0.03$ 0.684 ±$\pm 0.003$ 0.6921 ±$\pm 0.0054$ 0.2475 ±$\pm 0.0015$
Zinc. We experimented with the ZINC-12K molecular dataset (Sterling & Irwin, 2015; Gómez-Bombarelli et al., 2018), where we maintain a 500k parameter budget, in line with previous works.
Table 4: Test results on ZINC and MOLHIV. The first and second best results are color-coded.
Method ZINC ($\downarrow$) MOLHIV ($\uparrow$)
GCN 0.321 ±$\pm 0.009$ 76.06 ±$\pm 0.97$
GIN 0.163 ±$\pm 0.004$ 75.58 ±$\pm 1.40$
GSN 0.101 ±$\pm 0.010$ 80.39 ±$\pm 0.90$
CIN 0.079 ±$\pm 0.006$ 80.94 ±$\pm 0.57$
GPS 0.070 ±$\pm 0.004$ 78.80 ±$\pm 1.01$
GINE-MLP-Mixer 0.073 ±$\pm 0.001$ 79.97 ±$\pm 1.02$
GINE-ViT 0.085 ±$\pm 0.004$ 77.92 ±$\pm 1.42$
FULL 0.087 ±$\pm 0.003$ 76.54 ±$\pm 1.37$
OSAN 0.177 ±$\pm 0.016$
POLICY-LEARN (T = 2) 0.120 ±$\pm 0.003$ 79.13 ±$\pm 0.60$
RANDOM (T = 2) 0.136 ±$\pm 0.005$ 77.55 ±$\pm 1.24$
GIN+CSE 0.092 ±$\pm 0.002$ 77.44 ±$\pm 1.87$
HyMN (GIN, T=1) 0.080 ±$\pm 0.003$ 80.36 ±$\pm 1.23$
HyMN (GIN, T=2) 0.083 ±$\pm 0.002$ 81.01 ±$\pm 1.17$
We compared the test MAE using HyMN to other sampling approaches (Qian et al., 2022; Bevilacqua et al., 2024), a full-bag Subgraph GNN, a Graph Transformer (GPS) (Rampášek et al., 2022) , two expressive GNN baselines (GSN, CIN) (Bouritsas et al., 2022; Bodnar et al., 2021a) and the two alternative architectures Graph-ViT and Graph MLP-Mixers by (He et al., 2023). As can be seen from Table 4, our hybrid method can outperform a full-bag Subgraph GNN as well as previously proposed subsampling based approaches. Additionally, we perform competitively with CIN , which takes into account higher-order interactions and explicitly models ring-like structures. Non-message-passing based architectures can perform slightly better in absolute terms, but for higher run-time (see comparison with GPS in Appendix G). We highlight that additionally using subgraphs can outperform purely using the centrality-based encodings (Q4). We last refer to Appendix E for additional results on the Zinc and Peptides datasets obtained when sampling the bag based on centrality measures different than SC.
@@
Summary. In reference to the questions enlisted above, we conclude the following. (A1) The results from substructure counting and on the OGB benchmarks suggest a positive answer to Q1: SC-based sampling significantly outperformed random sampling on both, and matched the performance of the learnable POLICY-LEARN. (A2) HyMN was easily applied to the larger Peptides datasets, with strong empirical performance and favorable inference and training run-times. This indicates our approach broadens the applicability of Subgraph GNNs to larger graphs. (A3) Beyond Peptides, HyMN also attained remarkable performance on OGB benchmarks, and performed competitively on ZINC. This suggests a generally positive answer to Q3. (A4) We observe that CSEs can enhance the performance of standard MPNNs (see Table 3) and subsampled Subgraph GNNs (see, e.g., MOLHIV and MOLBACE in Table 2). We note, however, they were not beneficial on MOLTOX21 (Table 2).
6 Conclusion

### v2

@@ 4.1 Our Approach
Figure 2: An overview of the pipeline for HyMN. We calculate the CSE for each node, sample $T$ node-marked subgraphs with the highest Subgraph Centrality and concatenate the CSE with the initial node features.
@@
$C_{v}^{\mathsf{CSE}}=\left[1,\frac{(A)^{1}_{vv}}{1!},\frac{(A)^{2}_{vv}}{2!},\frac{(A)^{3}_{vv}}{3!},\dotsc,\frac{(A)^{K}_{vv}}{K!}\right].$ (4)
As $K\to\infty$, the sum of these terms clearly coincides with the SC of the node for the default $\beta=1$. We refer readers to Section F.5, for considerations on how CSEs compare with RWSEs.
@@
In view of (1) and (2), we dub our approach HyMN, as in Hybrid Marking Network. These steps are depicted in Figure 2 and described in Algorithm 1 (Appendix C). We note the following. First, for a large enough $K$, centrality values can be approximated by directly summing over the computed $K$-order CSEs. Second, node marking does not require any alteration of the original graph topology, making it unnecessary to store the subgraph connectivity. We thus opt not to materialize the bag of subgraphs: we only record marking information in the feature tensor and implement a custom message-function that processes it in an equivariant way. From an engineering standpoint, this allows for further memory-complexity enhancement w.r.t. generic Subgraph GNN approaches.
4.2 Expressivity of Subgraph GNNs with CSEs
@@
We answer these questions with graph separation arguments (Xu et al., 2018; Morris et al., 2019) and highlight how, in fact, the two approaches are not generally comparable. We demonstrate that subsampling Subgraph GNNs with high-centrality marking does not subsume CSE-enhanced MPNNs, while, at the same time, the discriminative power of the former approach is not fully captured by the latter. Since neither technique alone fully subsumes the other, our analysis emphasizes the advantages of combining them in HyMN for improved expressiveness. Proofs and additional details are reported in Appendix D.
@@
This result exposes a limitation of subsampled Subgraph GNNs in distinguishing between two non-isomorphic graphs with differing closed walks, features which are, instead, captured by CSEs. Notably, as discussed in Proposition D.5 (Appendix D), a full-bag approach is capable of capturing CSEs. This observation suggests that while CSEs do not universally enhance the expressiveness of any Subgraph GNN, they can be beneficial when subsampling a limited number of subgraphs.
Taken together, Theorems 4.1 and 4.2 indicate that leveraging SC both as a SE and as a means for subgraph sampling is advantageous in terms of discriminative power, justifying the integration of the two techniques in HyMN.
@@ 5 Experiments
  • (Q2) Can HyMN efficiently scale to graphs out of reach for Subgraph GNNs? How does it perform thereon?
@@
Figure 3: Comparing sampling approaches for counting triangles (performance metric is Mean Absolute Error, the lower the better). The average graph size is 59.33, corresponding to the average bag-size of a full-bag Subgraph GNN.
Synthetic experiment for counting substructures. The ability of a model to count local substructures is an acknowledged way of evaluating its expressive power (Bouritsas et al., 2022; Arvind et al., 2020; Tahmasebi et al., 2020). In order to answer (Q1) and test the efficacy of subgraph sampling with node centrality, we explored the ability of a Subgraph GNN to (learn to) count different small substructures as we increase the number of subgraphs in our bag. We closely followed the experimental procedure of (Chen et al., 2020), but modified the data generation process to render the task more challenging and informative [^5] In particular, we considered larger graphs with a similar number of nodes to correct an undesired correlation between the graph size and the task targets observed in the original data (see Section G.2).. We compared the performance of sampling subgraphs based on different approaches: random sampling and sampling based on the highest values of different centrality measures. No CSEs are employed in this setting. We remark that this experiment is intrinsically different than what explored in Section 3.2. There, we studied the correlation between subgraph counts and marking-induced perturbations on untrained models; here, we train models to regress these values (potentially making use of the additionally provided marking information). Figure 3 reports results for triangle counting, expressed in terms of Test Mean Absolute Error (MAE). These results demonstrate the significant improvement afforded by high-centrality sampling over random sampling with fewer subgraphs (Q1). Additionally, it shows that SC-based sampling generally outperformed other centrality measures, demonstrating the benefits of focusing on walk-based centralities, in alignment with our analysis in Section 3. The full set of results for additional substructures is reported in Appendix F.
@@
Table 2: Results on OGB datasets (Test ROC-AUC). The first and second best results for each task are color-coded. Symbol ‘-’ indicates the method is not (found to have been) benchmarked on the specific dataset.
Method MOLHIV MOLBACE MOLTOX21
GCN 76.06 ±$\pm 0.97$ 79.15 ±$\pm 1.44$ 75.29 ±$\pm 0.69$
GIN 75.58 ±$\pm 1.40$ 72.97 ±$\pm 4.00$ 74.91 ±$\pm 0.51$
FULL 76.54 ±$\pm 1.37$ 78.41 ±$\pm 1.94$ 76.25 ±$\pm 1.12$
OSAN (T = 10) - 76.30 ±$\pm 3.00$
Subsampled Subgraph GNNs MAG-GNN (T = 2) 77.12 ±$\pm 1.13$
RANDOM (T = 2) 77.55 ±$\pm 1.24$ 75.36 ±$\pm 4.28$ 76.65 ±$\pm 0.89$
CS-GNN (T = 2) 77.72 ±$\pm 0.76$ 80.58 ±$\pm 1.04$
POLICY-LEARN (T = 2) 79.13 ±$\pm 0.60$ 78.40 ±$\pm 2.85$ 77.47 ±$\pm 0.82$
HyMN (GIN, T=2) no CSE 79.77 ±$\pm 0.70$ 78.22 ±$\pm 4.02$ 77.68 ±$\pm 0.71$
@@
RANDOM (T=5) 77.30 ±$\pm 2.56$ 78.14 ±$\pm 2.36$ 76.62 ±$\pm 0.89$
CS-GNN (T = 5) 79.09 ±$\pm 0.90$ 79.64 ±$\pm 1.43$
POLICY-LEARN (T=5) 78.49 ±$\pm 1.01$ 78.39 ±$\pm 2.28$ 77.36 ±$\pm 0.60$
HyMN (GIN, T=5) no CSE 79.62 ±$\pm 1.14$ 78.57 ±$\pm 1.31$ 77.82 ±$\pm 0.59$
@@
Method Precomp. (s) Train (s/epoch) Test (s) Pept.-Func ($\uparrow$) Pept.-Struct ($\downarrow$)
@@
GPS 20.87 ±$\pm 0.43$ 8.39 ±$\pm 0.05$ 0.611 ±$\pm 0.005$ 0.6535 ±$\pm 0.0041$ 0.2500 ±$\pm 0.0005$
Graph-ViT 29.12 ±$\pm 0.61$ 6.78 ±$\pm 0.01$ 0.709 ±$\pm 0.009$ 0.6942 ±$\pm 0.0075$ 0.2449 ±$\pm 0.0016$
G-MLP-Mixer 29.52 ±$\pm 0.69$ 6.87 ±$\pm 0.03$ 0.684 ±$\pm 0.003$ 0.6921 ±$\pm 0.0054$ 0.2475 ±$\pm 0.0015$
Method Train (s/epoch) Test (s) MalNet-T. ($\uparrow$)
GIN 7.16 ±$\pm 0.04$ 0.70 ±$\pm 0.03$ 91.10 ±$\pm 0.98$
HyMN (T=1) no CSE 8.79 ±$\pm 0.01$ 0.94 ±$\pm 0.01$ 92.84 ±$\pm 0.52$
HyMN (T=1) 9.31 ±$\pm 0.02$ 1.02 ±$\pm 0.01$ 92.54 ±$\pm 0.75$
HyMN (T=2) no CSE 11.71 ±$\pm 0.17$ 1.38 ±$\pm 0.01$ 92.18 ±$\pm 0.64$
HyMN (T=2) 11.97 ±$\pm 0.06$ 1.40 ±$\pm 0.02$ 92.44 ±$\pm 0.35$
GPS (Perf.) 59.34 ±$\pm 0.52$ 6.43 ±$\pm 0.02$ 92.14 ±$\pm 0.24$
GPS (BigBird) 112.63 ±$\pm 1.26$ 17.65 ±$\pm 0.05$ 91.02 ±$\pm 0.48$
GPS (Transf.)* 162.10 ±$\pm 1.37$ 14.97 ±$\pm 0.04$ 90.85 ±$\pm 0.68$
Table 4: Results and timing comparisons using a GeForce RTX 2080 8 GB for the MalNet-Tiny dataset. The model marked with * did not fit in memory with batch size 16, and was trained with batch size 4. Both HyMN and GPS employ a GIN backbone.
OGB. We tested HyMN on several datasets for graph property prediction from the OGB benchmark (Hu et al., 2020b). We focus, in particular, on molecular graphs, tasked to predict global properties such as their toxicity or their ability to inhibit HIV replication. To further examine (Q1), Table 2 shows the performance of our approach in relation to MPNNs, a full-bag Subgraph GNN and other subgraph sampling policies with the same number of subgraphs. Notably, even without using the centrality-based encoding (HyMN w/out CSE), our method matches the performance of a learnable sampling policy (POLICY-LEARN (Bevilacqua et al., 2024)) and consistently outperforms MPNNs and random sampling policies. Additionally, we observe that augmenting node features with CSEs can significantly increase performance on MOLHIV and MOLBACE, outperforming a full-bag Subgraph GNN (Q4). These results suggest that centrality sampling is effective and that additionally incorporating centrality information can lead to performance improvements on real-world datasets, aligning with our findings in Section 3. We last highlight the exceptional result HyMN obtains, with $T=2$, on the MOLHIV benchmark ($81.01\pm 1.17$). This model outperforms strong GNN baselines, as reported in Section F.2, Table 8 (Q3). Run-time comparisons for this dataset are additionally enclosed in Appendix H.
Peptides. In order to evaluate the ability of HyMN to scale to larger graphs (Q2), we experimented on the Peptides datasets from the LRGB benchmark (Dwivedi et al., 2022), where the task is to predict global structural and functional properties of peptides, represented as graphs. The average number of nodes in these graphs is 150.94, so it is difficult for a full-bag Subgraph GNN to process. Additionally, using the centrality encoding to sample just one or two additional node-marked subgraphs can improve performance on both datasets. We also outperform GPS (Rampášek et al., 2022) (a Graph Transformer) and match the performance of Graph-MLP-Mixer (He et al., 2023) (Q3). Our timing experiments on Peptides-Func demonstrate that we are significantly more efficient than both of these approaches, only increasing the training time per epoch over a GCN by $42\%$ and the inference time by $25\%$ using ‘HyMN (GCN, T=1)’ (Q2).
MalNet-Tiny. This is a code dataset consisting of function call graphs with up to 5,000 nodes (Freitas et al., 2021). Models are asked to predict their malware class, or identify them as benign otherwise. As well as being from a significantly different domain, these graphs are considerably larger than the graph benchmarks considered above, allowing us to showcase HyMN’s ability to scale to even larger graphs (Q2). We compare to the GPS architecture (Rampášek et al., 2022), instantiated with various (sparse) variations of its Graph Transformer component and a GIN message-passing backbone, the same we employ for HyMN. The results are shown in Table 4. Whilst, full-bag Subgraph Graph Neural Networks cannot scale to this dataset, our method, HyMN, is able to outperform sparse Transformer-based approaches at a fraction of the runtime: our approach is, almost six times faster than the best GPS variant both in training and inference.
Reddit. To further showcase the ability of HyMN to scale to large graphs and to evaluate the effectiveness of the method with a different data modality, we experimented with the REDDIT-BINARY (RDT-B) dataset (Morris et al., 2020a). The task involves the classification of social networks that have a high average number of nodes $(429)$, so full-bag approaches cannot be applied. We compare our approach with a learnable sampling policy (POLICY-LEARN (Bevilacqua et al., 2024)) and to two expressive higher order GNNs (SIN and CIN) (Bodnar et al., 2021a, b). HyMN exhibits on-par performance with the best-performing method (5) which is a learnable policy, highlighting the effectiveness of our sampling approach.
Table 5: Results on RDT-B (Accuracy).
Method RDT-B
GIN 92.4 ±$\pm 2.5$
SIN 92.2 ±$\pm 1.0$
CIN 92.4 ±$\pm 2.1$
FULL OOM
RANDOM (T=20) 92.6 ±$\pm 1.5$
RANDOM (T=2) 92.4 ±$\pm 1.0$
POLICY-LEARN (T=2) 93.0 ±$\pm 0.9$
HyMN (T=2) 93.2 ±$\pm 2.2$
@@
Summary. In reference to the questions enlisted above, we conclude the following. (A1) The results from substructure counting and on the OGB benchmarks suggest a positive answer to Q1: SC-based sampling significantly outperformed random sampling on both, and matched the performance of the learnable POLICY-LEARN. (A2) HyMN was effortlessly applied to the larger Peptides and MalNet datasets, with strong empirical performance and extremely lightweight inference and training run-times. This demonstrates that effective subgraph sampling unlocks the applicability of Subgraph GNNs to larger graphs, even orders of magnitude larger than those full-bag methods could possibly be applied to. (A3) Beyond these two datasets, HyMN also attained remarkable performance on OGB benchmarks, and performed competitively on ZINC (see Section F.2). This suggests a positive answer to Q3. (A4) We observe that CSEs can enhance the performance of standard MPNNs (see Table 4) and subsampled Subgraph GNNs (see, e.g., MOLHIV and MOLBACE in Table 2). We note, however, they were not always beneficial (see MOLTOX21 Table 2 and MalNet Table 4).
6 Conclusions